OpenStudy (anonymous):

write a system of equations in two variables that has (5, -1) as its only solution

OpenStudy (aum):

x = 5, y = -1 Let us consider one equation as: x + y = 5 + (-1) = 4 Another equation could be: 2x + y = 2(5) + (-1) = 10 - 1 = 9 So the two equations are: x + y = 4 2x + y = 9

OpenStudy (aum):

That is just one set of equations. There are infinite such sets that will all yield (5,-1) as its only solution.

OpenStudy (anonymous):

Thanks

OpenStudy (aum):

You are welcome.